OpenStudy (trixitang):

Find the area of a circle with a diameter of 22 inches. Use 3.14 for pi. (1 point) 379.94 in2 452.16 in2 527.52 in2 621.72 in2

OpenStudy (trixitang):

im really bad at math, so what would i do ?

OpenStudy (er.mohd.amir):

radius=22/2=11 so area=pi*radius*radius

OpenStudy (er.mohd.amir):

area=3.14*11*11=?

OpenStudy (trixitang):

A

OpenStudy (trixitang):

379.94

OpenStudy (er.mohd.amir):

correct

OpenStudy (trixitang):

can you help with another ?

OpenStudy (trixitang):

A cone has a volume of 24 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of? (1 point) 8 in3 24 in3 48 in3 72 in3

OpenStudy (er.mohd.amir):

72 in3

OpenStudy (trixitang):

do you mind helping with more ?

OpenStudy (trixitang):

A cone has a volume of 12 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of? (1 point) 6 in3 24 in3 36 in3 48 in3

OpenStudy (er.mohd.amir):

36 in3

OpenStudy (trixitang):

I have like 5 more questions, is that okay

OpenStudy (trixitang):

A cone has a volume of 6 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of? (1 point) 12 in3 18 in3 24 in3 36 in3

OpenStudy (er.mohd.amir):

Q are same ans=volume of cone *3

OpenStudy (trixitang):

so i just multiply whatever the volume is by 3 ?

OpenStudy (er.mohd.amir):

no if vol of cone given and find vol of cylinder

OpenStudy (trixitang):

so for that one i put 18 in3 and for this one A cone has a volume of 5 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of? (1 point) 15 in3 20 in3 25 in3 30 in3 i put the first one, i think im wrong

OpenStudy (er.mohd.amir):

wait just a min

OpenStudy (er.mohd.amir):

vol of cylinder =3 * vol of cone use this

OpenStudy (er.mohd.amir):

yes last ans is 15 in3 is correct

OpenStudy (mathmate):

@Er.Mohd.AMIR http://prntscr.com/bbibl0